Given the ellipse with equation:(x-5)^2/16+(y=1)^2/25=1
Find the center, the length of the minor axis, the length of the major axis, and tell if it is oriented left/right or up/down.

Answer by stanbon(75887)

Given the ellipse with equation:(x-5)^2/16+(y+1)^2/25=1
Find the center: (5,-1)
----------------------------------
the length of the minor axis:: 2*sqrt(16) = 2*4 = 8
-------------------------------------------------
the length of the major axis:: 2*sqrt(25) = 2*5 = 10
-------------------------------------------------
and tell if it is oriented left/right or up/down: up/down.
-----------------------
Cheers,
Stan H.
================